# Swing: Short-cutting Rings for Higher Bandwidth Allreduce

### 17 Jan 2024 v1

@@ Abstract
The allreduce collective operation accounts for a significant fraction of the runtime of workloads running on distributed systems. One factor determining its performance is the distance between communicating nodes, especially on networks like torus, where a higher distance implies multiple messages being forwarded on the same link, thus reducing the allreduce bandwidth. Torus networks are widely used on systems optimized for machine learning workloads (e.g., Google TPUs and Amazon Trainium devices), as well as on some of the Top500 supercomputers. To improve allreduce performance on torus networks we introduce Swing, a new algorithm that keeps a low distance between communicating nodes by swinging between torus directions. Our analysis and experimental evaluation show that Swing outperforms by up to $3x$ existing allreduce algorithms for vectors ranging from 32B to 128MiB, on different types of torus and torus-like topologies, regardless of their shape and size.
@@ 1 Introduction and Motivation
Allreduce is a collective operation used to aggregate vectors among a set of nodes and to distribute back to them the aggregated result. Among others, allreduce is widely used to perform distributed gradient aggregation during the training of deep learning models [10]. Several studies have shown that it can account for up to 40% of the total training time [44, 28, 39] and between 19% and 30% of the total core hours in MPI jobs running on production supercomputers [14].
Researchers proposed several allreduce algorithms [43, 6, 25], and the most performing one depends on a combination of vector size, number of nodes, and physical topology [42, 23]. Those algorithms perform a predefined number of steps and, at each step, each node sends and receives data to and from some predetermined nodes. Different trade-offs exist between the number of steps to perform (more critical for allreduce on small vectors) and the total number of bytes it transmits (more relevant for larger allreduce). However, a third factor that must be considered when designing a new collective algorithm is the physical distance between communicating nodes [43, 33, 38, 42].
This is particularly relevant on networks that do not provide full bisection bandwidth such as torus, since the higher the distance, the higher the number of flows sharing the same links. Torus networks are widely used, both on systems optimized for running machine learning (ML) workloads (e.g., Google deploys TPUs v4 on a 3D torus-like network [27], and Amazon deploys Trainium devices on a 2D torus [7]), and on some of the top supercomputers [2, 11] (e.g., Fugaku uses a 6D torus [4]).
@@
We show the importance of the distance between communicating nodes in the allreduce through an example. In Fig. 1, we show a $16$ node 1D torus (we only show a subset of the nodes since the communications are symmetric). We assume minimal (i.e., shortest path) routing and we show the communications performed by the bandwidth-optimal recursive doubling algorithm [43] (also known as Rabenseifner algorithm [35], which we describe more in detail in Sec. 2.3.3), and by the Swing algorithm (that we propose in this work).
@@
It is thus clear how even if two different algorithms transmit the same number of bytes and perform the same number of steps, they might be practically characterized by different performance, depending on the network characteristics and the distance between communicating nodes. In this example, we have shown an extreme case using a 1D torus. However, similar effects can happen on any topology that does not provide full bisection bandwidth.
@@
For these reasons, this work makes the following contributions:
@@ 2.2 Notation and Model
$\delta(s)$ Distance between $2$ communicating nodes at step $s$
@@ 2.3.5 Other Approaches
Some approaches use linear programming formulations for finding the best collective algorithm given a network specification and the size of the collective [41, 13]. However, this requires solving an NP-hard problem that grows exponentially with scale. Finding a solution for 128 nodes requires up to 11 hours [41], and a new solution might need to be found when changing the number of nodes or the size of the collective.
@@ 3.1.1 Bandwidth-optimal Algorithm
Where $\rho(s)=\sum_{i=0}^{s}-2^{i}=\frac{1-(-2)^{s+1}}{3}$. This selection of the communicating peer leads to a communication pattern like the one shown in Figure 1 for a $16$ nodes 1D torus. We observe how, at each step, the communicating peer of each node swings from left to right and vice versa (hence the algorithm’s name). Intuitively, unlike recursive doubling, each node reaches distant nodes in fewer hops by short-cutting the ring.
@@
%**** main.tex Line 400 ****for s in range(step, int(log2(p))):
@@ 5.2 Performance on Rectangular Tori
On the other hand, the shape of the torus negatively impacts the bucket algorithm’s latency deficiency. Indeed, if some of the torus dimensions are larger than the others, some of the $2D$ concurrent collectives might move from dimension $i$ to dimension $i+1$, whereas there are still collectives running on dimension $i+1$. We show this through an example in Fig. 10, where at step 1 one bidirectional ring moves from the vertical to the horizontal dimension while there is still the other bidirectional ring running.
Figure 10: First two steps of the Bucket algorithm on a 2x4 torus.
Figure 11: Goodput on 2D torus with $1024$ nodes and different rectangular shapes: (a) 64x16 (b) 32x8 (c) 256x4.
Figure 12: Goodput on higher-dimensional torus networks: (a) 2D 8x8 (b) 3D (8x8x8) (c) 4D (8x8x8x8).
@@ 5.3 Performance for 3D and 4D Torus
As discussed in Sec. 4 and summarized in Table 2, the performance of the allreduce algorithm for multidimensional torus also depends on the number of dimensions. Thus, we evaluate the performance of the different allreduce algorithms on $8^{2}$, $8^{3}$, and $8^{4}$ torus networks. As discussed in Sec. 2.3.1, the Hamiltonian ring algorithm only works for 2D torus networks, and we do not include it in the results for 3D and 4D torus.
When increasing the number of dimensions, the goodput gain of Swing increases because, as shown in Table 2 and discussed in Sec. 4, the congestion deficiency drops to $3\%$ on 3D torus and to $0.8\%$ on 4D torus. Consequently, for 3D and 4D torus networks, Swing outperforms by up to 2x all existing algorithms on allreduce ranging from 32B to 2GB.
@@ 5.4.1 Performance on HammingMesh
We show in Fig. 13 the performance of the different algorithms for a Hx2Mesh network with $4096$ nodes (2x2 boards arranged in a 32x32 configuration). For such configuration, Swing outperforms the state-of-the-art algorithms at any size, up to 2.5x for 2MiB allreduce. Moreover, because of the lower congestion deficiency, we observe how the peak Swing performance is higher compared to a 2D torus with the same number of nodes (Fig. 7). Last, we also observe a runtime reduction for all the algorithms for small allreduce, since nodes on the same board on HammingMesh are connected through PCB traces and thus have lower latency compared to optical network cables.
@@ 5.5 Summary
First, by observing the performance on square torus networks, we see that Swing median and maximum goodput gain increases with the network size (going from 8x8 to 128x128 torus). As discussed in Sec. 5.1, the maximum negative gain also increases with network size but, however, is at most around $20\%$, and only occurs for large allreduce ($\geq$ 128MiB).
@@ 6 Discussion
In all our experiments we used minimal adaptive routing. I.e., packets are forwarded on the least congested shortest path. Because in all the analyzed algorithms, each node only communicates with others on the same row/column, traffic is evenly distributed across links. Considering also non-shortest paths would unnecessarily increase network traffic and decrease performance for all the algorithms.
@@
On full-bandwidth topologies (e.g., non-blocking fat tree), both Swing and recursive doubling will not have any congestion deficiency, and they will be characterized by the same performance.
@@ A.1 Power of Two Number of Nodes
From Lemma A.4, we know that both sides are in the range $(-p,p)$. Thus, the two sides are congruent only if: i) they have the same sign and are equal, or; ii) they have different signs, and by summing $p$ on the negative side, we get the positive side. Since each side is the sum of distinct powers of $-2$, case i) is only possible if the two sequences of steps are equal. To prove that case ii) is impossible, let us consider the case where the left side is negative (the other case is analogous). Because $p=2^{a}$ for some $a\in\mathbb{N}$, and because $2^{a}=(-2)^{a+1}+(-2)^{a}$, Eq. 5 becomes:
$\begin{split}&\sum_{i=0}^{s_{0}}(-2)^{i}+\ldots+\sum_{\mathclap{i=s_{k-2}+1}}^{s_{k-1}}(-2)^{i}+(-2)^{a+1}+(-2)^{a}=\\ &=\sum_{i=0}^{t_{0}}(-2)^{i}+\ldots+\sum_{\mathclap{i=t_{k-2}+1}}^{t_{h-1}}(-2)^{i}\end{split}$
However, because both sides are sums of distinct powers of $-2$, they can be equal only if the two sequences of steps are equal, which implies that there must be a $t_{i}$ such that $t_{i}=a+1$. Because steps can be at most equal to $\log_{2}(p)-1=a-1$, this is not possible.
@@ A.2 Non-Power of Two Number of Nodes
The correctness proof in Sec. A.1 assumes $p$ is a power of $2$ (needed by Lemma A.4). If $p$ is not a power of $2$, Lemma A.4 only holds until the second-last step. If $p^{\prime}$ is the largest power of $2$ smaller than $p$, in the second last step, the data sent by each node reached $p^{\prime}-1$ nodes. Thus, we need to guarantee that in the last step: i) no nodes receive data it already received; ii) each node reaches the remaining $p-p^{\prime}-1$ nodes. To guarantee property i), it is enough for each node to pre-compute the blocks $b_{i}$ it will send at each step and if it would send a block twice, send that only in the last step.

### 04 Mar 2024 v2

@@ Abstract
The allreduce collective operation accounts for a significant fraction of the runtime of workloads running on distributed systems. One factor determining its performance is the number of hops between communicating nodes, especially on networks like torus, where a higher number of hops implies multiple messages being forwarded on the same link, thus reducing the allreduce bandwidth. Torus networks are widely used on systems optimized for machine learning workloads (e.g., Google TPUs and Amazon Trainium devices), as well as on some of the Top500 supercomputers. To improve allreduce performance on torus networks we introduce Swing, a new algorithm that reduces the number of hops between communicating nodes by swinging between torus directions. Our analysis and experimental evaluation show that Swing outperforms by up to $3x$ existing allreduce algorithms for vectors ranging from 32B to 128MiB, on different types of torus and torus-like topologies, regardless of their shape and size.
@@ 1 Introduction and Motivation
Allreduce is a collective operation used to aggregate vectors among a set of nodes and to distribute the aggregated result back to them. Allreduce is widely used to perform distributed gradient aggregation when training deep learning models [10]. Studies have shown that it can account for up to 40% of the total training time [51, 32, 45] and between 19% and 30% of the total core hours in MPI jobs running on production supercomputers [14].
In this work, we design a new allreduce algorithm optimized for torus-like networks. Torus networks are widely used, both on systems optimized for running machine learning (ML) workloads and on some of the top supercomputers [2, 11] (e.g., Fugaku uses a 6D torus [4]). Although torus networks are characterized by a lower bisection and global bandwidth compared to other topologies (e.g., Clos), their simplicity and lower cost allow running some workloads such as ML training in a more cost-effective way, since their communication is often arranged as a 3D logical torus [26].
Seen from a different perspective, torus networks trade off a lower cost for a lower bisection bandwidth, which, however, is enough to train most ML models efficiently [26]. This is the reason why many systems optimized for ML training rely on torus-like topologies. These include, for example, Google TPUs [31] (a TPU v5p pod connects $9000$$ chips on a 3D torus [15]), Amazon Trainium devices [7] (16 chip on a 2D torus), Graphcore IPU-POD [24] (connecting 64 chips on a 2D torus), Enflame [40] (2D torus).
Researchers proposed several allreduce algorithms [50, 6, 29], and the most performing one depends on a combination of vector size, number of nodes, and physical topology [49, 27]. Those algorithms perform a predefined number of steps and, at each step, each node sends and receives data to and from some predetermined nodes. Different trade-offs exist between the number of steps to perform (more critical for allreduce on small vectors) and the total number of bytes it transmits (more relevant for larger allreduce). However, a third factor that must be considered when designing a new collective algorithm is the number of hops between communicating nodes [50, 38, 44, 49]. This is particularly relevant on networks that do not provide full bisection bandwidth such as torus, since the higher the number of hops, the higher the number of flows sharing the same links.
@@
We show the importance of the number of hops in the allreduce through an example. In Fig. 1, we show a $16$ node 1D torus (we only show a subset of the nodes since the communications are symmetric). We assume minimal (i.e., shortest path) routing and we show the communications performed by the bandwidth-optimal recursive doubling algorithm [50] (also known as Rabenseifner algorithm [41], which we describe more in detail in Sec. 2.3.3), and by the Swing algorithm (that we propose in this work).
@@
It is thus clear how even if two different algorithms transmit the same number of bytes and perform the same number of steps, they might have different performance, depending on the network characteristics and the distance between communicating nodes. In this example, we have shown an extreme case using a 1D torus. However, similar effects can happen on any topology that does not provide full bisection bandwidth.
@@
This work makes the following contributions:
@@ 2.2 Notation and Model
$\delta(s)$ Number of hops between $2$ communicating nodes at step $s$
@@ 2.3.5 Other Approaches
Automatic Generation of Collective Algorithms
Some approaches use linear programming formulations for finding the best collective algorithm given a network specification and the size of the collective [48, 13]. However, this requires solving an NP-hard problem that grows exponentially with scale. Finding a solution for 128 nodes requires up to 11 hours [48], and a new solution might need to be found when changing the number of nodes or the size of the collective. This makes generating collective algorithms for large systems like the $9000$ nodes Google’s TPU v5p pod [15] challenging or even impossible. On the contrary, Swing can seamlessly run on any number of nodes. Moreover, unlike Swing, some of these solutions do not explicitly model the congestion deficiency.
Topology Reconfiguration and In-Network Compute
Other solutions improve allreduce performance by re-configuring the network topology according to the specific traffic pattern [51]. Swing is orthogonal to these approaches and, by reducing the network congestion, can make the expensive network re-configurations less frequent. Last, some algorithms exploit in-network compute capabilities of programmable switches [46, 34, 21, 17] to aggregate data directly in the network, reducing network traffic and improving performance. However, unlike Swing, these solutions require specific switches to be deployed in the network, whereas Swing can seamlessly run on any network.
@@ 3.1.1 Bandwidth-optimal Algorithm
Where $\rho(s)=\sum_{i=0}^{s}(-2)^{i}=\frac{1-(-2)^{s+1}}{3}$. This selection of the communicating peer leads to a communication pattern like the one shown in Figure 1 for a $16$ nodes 1D torus. We observe how, at each step, the communicating peer of each node swings from left to right and vice versa (hence the algorithm’s name). Intuitively, unlike recursive doubling, each node reaches distant nodes in fewer hops by short-cutting the ring.
@@
for s in range(step, int(log2(p))):
@@ 5.2 Performance on Rectangular Tori
On the other hand, the shape of the torus negatively impacts the bucket algorithm’s latency deficiency. Indeed, if some torus dimensions are larger than the others, some of the $2D$ concurrent collectives might move from dimension $i$ to dimension $i+1$, whereas there are still collectives running on dimension $i+1$. We show this through an example in Fig. 10, where at step $1$, one bidirectional ring moves from the vertical to the horizontal dimension while the other bidirectional ring is still running.
Figure 10: First steps of the Bucket algorithm on a 2x4 torus.
@@
Figure 11: Goodput on 2D torus with $1024$ nodes and different rectangular shapes: 64x16, 32x8, and 256x4.
@@ 5.3 Performance for 3D and 4D Torus
As discussed in Sec. 4 and summarized in Table 2, the performance of the allreduce algorithm for multidimensional torus also depends on the number of dimensions. Thus, we evaluate the performance of the different allreduce algorithms on $8^{2}$, $8^{3}$, and $8^{4}$ torus networks.
Figure 12: Goodput on higher-dimensional torus networks: 2D 8x8, 3D (8x8x8), and 4D (8x8x8x8).
We report the evaluation result in Fig. 12. We do not include the Hamiltonian ring algorithm in the 3D and 4D torus results since it only works for 2D torus networks. When increasing the number of dimensions, the goodput gain of Swing increases because, as shown in Table 2 and discussed in Sec. 4, the congestion deficiency drops to $3\%$ on 3D torus and to $0.8\%$ on 4D torus. Consequently, for 3D and 4D torus networks, Swing outperforms by up to 2x all existing algorithms on allreduce ranging from 32B to 2GB.
@@ 5.4.1 Performance on HammingMesh
We show in Fig. 13 the performance of the different algorithms for a Hx2Mesh network with $4096$ nodes (2x2 boards arranged in a 32x32 configuration). For such configuration, Swing outperforms the state-of-the-art algorithms at any size, up to 2.5x for 2MiB allreduce. Moreover, because of the lower congestion deficiency, we observe how the peak Swing performance is higher compared to a 2D torus with the same number of nodes (Fig. 7). Last, we also observe a runtime reduction for all the algorithms for small vectors, since nodes on the same board on HammingMesh are connected through PCB traces, with lower latency than optical network cables.
@@ 5.5 Summary
First, by observing the performance on square torus networks, we see that Swing median and maximum goodput gain increases with the network size (going from 8x8 to 128x128 torus). As discussed in Sec. 5.1, the maximum negative gain also increases with network size, but is at most around $20\%$ and only occurs for large allreduce ($\geq$ 128MiB).
@@ 6 Discussion
In all our experiments, we used minimal adaptive routing. I.e., packets are forwarded on the least congested shortest path. Because in all the analyzed algorithms, each node only communicates with others on the same row/column, traffic is evenly distributed across links. Sending packets on non-shortest paths would unnecessarily increase network traffic and decrease performance.
@@
On full-bandwidth topologies (e.g., non-blocking fat trees), both Swing and recursive doubling will not have any congestion deficiency, and we expect them to have the same performance.
@@ 7 Conclusions
Acknowledgments
We thank the anonymous NSDI reviewers and our shepherd, Philip Levis, for their constructive feedback. This work was supported by Sapienza University under the SEED-2022 and "Progetti Grandi 2023" funding schemes; by UrbanTwin, an ETH Board Joint Initiatives project; and by EuroHPC-JU under grant agreement RED-SEA, No 055776. Daniele De Sensi is a member of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM).
@@ A.1 Power of Two Number of Nodes
From Lemma A.4, we know that both sides are in the range $(-p,p)$. Thus, the two sides are congruent only if: i) they have the same sign and are equal, or; ii) they have different signs, and by summing $p$ on the negative side, we get the positive side. Since each side is the sum of distinct powers of $-2$, case i) is only possible if the two sequences of steps are equal. To prove that case ii) is impossible, let us consider the case where the left side is negative (the other case is analogous). Because $p=2^{a}$ for some $a\in\mathbb{N}$, and because $2^{a}=(-2)^{a}$ (if $a$ is even [^2] If $a$ is odd, $p=2^{a}=(-2)^{a+1}+(-2)^{a}$, and the same considerations still hold.), Eq. 5 becomes:
$\begin{split}&\sum_{i=0}^{s_{0}}(-2)^{i}+\ldots+\sum_{\mathclap{i=s_{k-2}+1}}^{s_{k-1}}(-2)^{i}+(-2)^{a}=\\ &=\sum_{i=0}^{t_{0}}(-2)^{i}+\ldots+\sum_{\mathclap{i=t_{k-2}+1}}^{t_{h-1}}(-2)^{i}\end{split}$
However, because both sides are sums of distinct powers of $-2$, they can be equal only if the two sequences of steps are equal, which implies that there must be a $t_{i}$ such that $t_{i}=a$. This is impossible because the number of steps can be at most equal to $\log_{2}(p)-1=a-1$.
@@ A.2 Non-Power of Two Number of Nodes
The correctness proof in Sec. A.1 assumes $p$ is a power of $2$ (needed by the last part of Theorem A.5). If $p$ is not a power of $2$, the theorem only holds until the second-last step. If $p^{\prime}$ is the largest power of $2$ smaller than $p$, in the second last step, the data sent by each node reached $p^{\prime}-1$ nodes. Thus, we need to guarantee that in the last step: i) no nodes receive data it already received; ii) each node reaches the remaining $p-p^{\prime}-1$ nodes. To guarantee property i), it is enough for each node to pre-compute the blocks $b_{i}$ it will send at each step and if it would send a block twice, send that only in the last step.